The sum of two numbers is 50 and the sum of their squares is 2050 .what are the numbers?

Answer by shahbazn(6)

Let one number be x. other number = (50-x)
x^2=(50-x)^2 = 2050
x^2+2500-100x+x^2 = 2050
2x^2-100x+2500-2050=0
2x2-100x+450=0
x^2-50x+225=0 225=5*5*3*3
(x-5)(x-450=0 225=5*45
x=5 or 45
smaller number = 5
larger number =45